OpenStudy (anonymous):

add and simplify, if possible x^2+13x/x^2-7x + x^2-8x/x^2-7x

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2+13x}{x^2-7x} + \frac{x^2-8x}{x^2-7x}\] \[\Large \frac{x^2+13x+x^2-8x}{x^2-7x}\] \[\Large \frac{2x^2+5x}{x^2-7x}\] So \[\Large \frac{x^2+13x}{x^2-7x} + \frac{x^2-8x}{x^2-7x}=\frac{2x^2+5x}{x^2-7x}\]

OpenStudy (anonymous):

x(2x+5)/x(x-7)=(2x+5)/(x-7)